OpenStudy (anonymous):

Five cards are drawn from an ordinary deck of 52 playing cards. What is the probability of getting a full house? (3 of a kind, and 2 of a kind)

OpenStudy (saifoo.khan):

LET US BE FIRST CLEAR OF WHAT WE MEAN BY ONE KIND.... I TAKE IT AS SAME SUIT..THAT IS TWO SPADES AND THREE DIAMONDS SAY..OR LIKE THAT...ON THIS BASIS TOTAL NUMBER OF COMBINATIONS TO DRAW 5 CARDS OUT OF 52 IS 52C5............I THERE CARE 4 SUITS.WE WANT CARDS OF ANY 2 SUITS ONLY..SO LET US FIRST SELECT 2 SUITS FROM THE 4 AVAILBLE. THIS CAN BE DONE IN 4C2 WAYS=6 WAYS.................II NOW ONE COULD BE 2 CARDER AND ANOTHER COULD BE 3 CARDE SO THE NUMBER OF POSSIBILITIES ARE =2 FROM THE ABOVE 2 SELECTED SUITS......III THERE ARE 13 CARDS OF EACH SUIT. NUMBER OF WAYS OF GETTING 2 CARDS OF SAME SUIT IS 13C2.....IV NUMBER OF WAYS OF GETTING 3 CARDS OF SAME SUIT IS 13C3.....V SO TOTAL NUMBER OF COMBINATIONS TO GET FULL HOUSE ARE II*III*IV*V SUCCESSFULL COMBINATIONS =6*2*13C2*13C3 PROBABILITY=6*2*13C2*13C3/52C5

OpenStudy (anonymous):

Thank you so much for your help! :)

OpenStudy (saifoo.khan):

\[\Huge \color\red {U.r.WELCOME!.. :)}\]

OpenStudy (saifoo.khan):

Can i ask wht does "5369855" stands for?

OpenStudy (anonymous):

it's my name in letters :) like ABC = 1

OpenStudy (saifoo.khan):

OIC, Cool. Hard to guess. lol.

OpenStudy (anonymous):

Yep :) No one understands it haha

OpenStudy (saifoo.khan):

Lol.

OpenStudy (anonymous):

:D